You need 1275 mL of a 15% alcohol solution. On hand, you have a 10% alcohol mixture and a 95% alcohol mixture. How much of each mixture will you need to obtain the desired solution?

Answer by TimothyLamb(4379)

x = ml of 10% mix
y = ml of 95% mix
---
x + y = 1275
0.10x + 0.95y = 0.15*1275
---
put the system of linear equations into standard form
---
x + y = 1275
0.10x + 0.95y = 191.25
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = 1200
y = 75
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---